# ACEBench: Who Wins the Match Point in Tool Usage?

### v3

@@ C.1 Normal Examples
  "description": "Assess potential security threats and provide mitigation suggestions",
@@
  "description": "Manage user’s privacy settings",
@@
Figure 12: Example of Normal Single Turn Data.
@@
Figure 13: Example of Normal Multi Turn Data.
@@
Figure 14: Example of Normal Preference Data.
@@
  "description": "Examines baby’s vein visibility and provides a  recommendation to see a doctor.
@@
Figure 15: Example of Similar API Data.
@@
Figure 16: Example of Atom (enum) Data.
C.2 Special Example
@@
Figure 17: Example of Special Lose Param Data.
@@
Can you retrieve the tennis team details for the team named ’Team@1234’?”
@@
Figure 18: Example of Special Error Param Data.
@@
Figure 19: Example of Special Exceed Function Ability Data.
C.3 Agent Example
Figure 20: Example of Agent Data.
Figure 21: Example of Agent Data.
Appendix D Evaluation Inference Prompts
D.1 Normal Prompt
@@
You are an AI assistant with the role name "assistant." Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], replacing ApiName with the actual API name, key1, key2, etc., with the actual parameter names, and value1, value2, etc., with the actual parameter values. The output should start with a square bracket "[" and end with a square bracket "]". If there are multiple API requests, separate them with commas, for example:[ApiName(key1=’value1’,key2=’value2’,…),ApiName(key1=’value1’,key2=’value2’, …), …]. Do not include any other explanations, prompts, or API call results in the output. If the API parameter description does not specify otherwise, the parameter is optional (parameters mentioned in the user input need to be included in the output; if not mentioned, they do not need to be included). If the API parameter description does not specify the required format for the value, use the user’s original text for the parameter value. If the API requires no parameters, output the API request directly in the format [ApiName()], and do not invent any nonexistent parameter names.
@@
Figure 22: System Prompt for Nomal (Except Prefernce).
@@
You are an AI assistant, and your role is called assistant. Based on the given API description, dialogue history 1..t, and character profile, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], where ApiName is replaced with the actual API name, and key1, key2, etc., are replaced with the actual parameter names, and value1, value2 are replaced with the actual parameter values. The output should start with a "[" and end with a "]". If there are multiple API requests, they should be separated by commas, e.g., [ApiName(key1=’value1’, key2=’value2’, …), ApiName(key1=’value1’, key2=’value2’, …), …]. Do not output any other explanations, hints, or results of the API calls in the output. If the API parameter description does not specify special instructions, the parameter is optional (parameters mentioned in the user input or character profile should be included in the output, and if not mentioned, they should not be included). If the API parameter description does not specify the format for the parameter value, the parameter value should be taken from the user’s original text or character profile. If the API requires no parameters, the API request should be output as [ApiName()], with no fabricated parameter names.
@@
Figure 23: System Prompt for Nomal (Prefernce).
D.2 Special Prompt
@@
You are an AI system with the role name "system." Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the system should call in step t+1. Below are two specific scenarios:
@@
Figure 24: System Prompt for Special.
D.3 Agent Prompt
@@
Figure 25: System Prompt for Agent.
Appendix E Error Examples
E.1 Error Examples of Normal Data
{mdframed}
{"Escher_Artwork_Analysis_Tool": {
    "artwork_analysis": [
      {"artwork_id": "Waterfall",
       "analysis_type": "Symmetry Analysis"}]}}
(model output)
{“FontAdvisor_recommendDisplayFonts”:{……}}
Figure 26: Error Example of Noraml Wrong Function.
{mdframed}
(model output)
Figure 27: Error Example of Noraml Param Type.
{mdframed}
{ "SignalAccuracyEvaluator":{ "device_id":"GPS-7526“ }}
(model output)
{"SignalAccuracyEvaluator":{ "device_id":"GPS-7524“ }}
Figure 28: Error Example of Noraml Param Value.
{mdframed}
{"EcoDiversityMonitor_assessVegetation": {
    "area_identifier": "Amazon rainforest",
    "detailed_report": true}}
(model output)
Figure 29: Error Example of Wrong Output Format.
E.2 Error Examples of Special Data
{mdframed}
(model output)
  "Get_Middle_East_Cities": {
    "sort": "desc",
    "sort_by": "internet_speed",
    "size": 10,
    "page": 1
(The question did not mention the sort method)
Figure 30: Error Example of Special Error Blindness.
{mdframed}
(model output)
(Find missing necessary parameters, but not correctly identified)
Figure 31: Error Example of Special Vague Error Reporting.
Appendix F Formula for Overall Accuracy
The formula for calculating the Overall Accuracy can be expressed as:
where the coefficients $A$, $B$, and $C$ are defined as:
F.1 Detailed Results
@@
Atom Single Turn Multi Turn Similar API Preference Summary
@@
Atom Single Turn Multi Turn Similar API Preference Summary

### v4

@@ C.1 Normal Examples
Single-Turn. The example of Normal Single-Turn Data is shown in Figure 14.
Multi-Turn. The example of Normal Multi-Turn Data is shown in Figure 15.
Preference. The example of Normal Preference Data is shown in Figure 16.
Similar APIs. The example of Normal Similar APIs Data is shown in Figure 17.
Atom. The example of Normal Atom Data is shown in Figure 18.
C.2 Special Examples
Incomplete. The example of Special Incomplete Data is shown in Figure 19.
Error. The example of Special Error Data is shown in Figure 20.
Irrelevant. The example of Special Irrelevant Data is shown in Figure 21.
C.3 Agent Examples
The example of Agent Data is shown in Figure 22 and Figure 23.
Appendix D Evaluation Inference Prompts
D.1 Normal Prompt
The main evaluation inference prompt for Normal Data is shown in Figure 24 and Figure 25.
D.2 Special Prompt
The main evaluation inference prompt for Special Data is shown in Figure 26.
D.3 Agent Prompt
An example of the evaluation inference prompt for Agent Data in a specific scenario is shown in Figure 27.
Appendix E Formula for Overall Accuracy
The formula for calculating the Overall Accuracy can be expressed as:
where the coefficients $A$, $B$, and $C$ are defined as:
Appendix F Detailed Results
The accuracy evaluation of different models on English Data is shown in Table 6, and the accuracy evaluation of different models on Chinese Data is shown in Table 7.
Appendix G Error Examples
G.1 Error Examples of Normal Data
Wrong Function Name. An example of wrong function name is shown in Figure 28.
Wrong Param Type. An example of wrong param type is shown in Figure 29.
Wrong Param Value. An example of wrong param value is shown in Figure 30.
Wrong Output Format. An example of wrong output format is shown in Figure 31.
G.2 Error Examples of Special Data
Error Detection Figure 32 shows an example of error detection.
Error Correction Figure 33 shows an example of error correction.
@@ G.2 Error Examples of Special Data
        "description": "Assess potential security
        threats and provide mitigation suggestions",
@@
        "description": "Manage user’s privacy",
@@
Figure 14: An example of Normal Single-Turn Data.
@@
Figure 15: An example of Normal Multi-Turn Data.
@@
Figure 16: An example of Normal Preference Data.
@@
  "description": "Examines baby’s vein visibility and recommends seeing a doctor.
@@
Figure 17: An example of Similar API Data.
@@
Figure 18: An example of Atom (enum) Data.
@@
Figure 19: An example of Special Incomplete Data.
@@
Can you retrieve the tennis team details named ’Team@1234’?”
@@
Figure 20: An example of Special Error Param Data.
@@
Figure 21: An example of Special Irrelevant Data.
Figure 22: An example of Agent Data: A partially correct representation.
Figure 23: An example of Agent Data: A fully accurate representation.
@@
You are an AI assistant with the role name "assistant". Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], replacing ApiName with the actual API name, key1, key2, etc., with the actual parameter names, and value1, value2, etc., with the actual parameter values. The output should start with a square bracket "[" and end with a square bracket "]". If there are multiple API requests, separate them with commas, for example:[ApiName(key1=’value1’,key2=’value2’,…),ApiName(key1=’value1’,key2=’value2’, …), …]. Do not include any other explanations, prompts, or API call results in the output. If the API parameter description does not specify otherwise, the parameter is optional (parameters mentioned in the user input need to be included in the output; if not mentioned, they do not need to be included). If the API parameter description does not specify the required format for the value, use the user’s original text for the parameter value. If the API requires no parameters, output the API request directly in the format [ApiName()], and do not invent any nonexistent parameter names.
@@
Figure 24: The inference prompt for Nomal (except Prefernce) Data.
@@
You are an AI assistant, and your role is called assistant. Based on the given API description, dialogue history 1..t, and character profile, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], where ApiName is replaced with the actual API name, and key1, key2, etc., are replaced with the actual parameter names, and value1, value2 are replaced with the actual parameter values. The output should start with a "[" and end with a "]". If there are multiple API requests, they should be separated by commas, e.g., [ApiName(key1=’value1’, key2=’value2’, …), ApiName(key1=’value1’, key2=’value2’, …), …]. Do not output any other explanations, hints, or results of the API calls in the output. If the API parameter description does not specify special instructions, the parameter is optional (parameters mentioned in the user input or character profile should be included in the output, and if not mentioned, they should not be included). If the API parameter description does not specify the format for the parameter value, it should be taken from the user’s original text or character profile. If the API requires no parameters, the API request should be output as [ApiName()], with no fabricated parameter names.
@@
Figure 25: The inference prompt for Nomal (Prefernce) Data.
@@
You are an AI system with the role name "assistant". Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the system should call in step t+1. Below are two specific scenarios:
@@
Figure 26: The inference prompt for Special Data.
@@
Figure 27: The inference prompt for Agent Data in a specific scenario.
@@
Atom Single-Turn Multi-Turn Similar API Preference Summary
@@
Atom Single-Turn Multi-Turn Similar API Preference Summary
@@
{mdframed}
{"Escher_Artwork_Analysis_Tool": {
    "artwork_analysis": [
      {"artwork_id": "Waterfall",
       "analysis_type": "Symmetry Analysis"}]}}
(model output)
{“FontAdvisor_recommendDisplayFonts”:{……}}
Figure 28: An error example of Wrong Function Name on Normal Data.
{mdframed}
(model output)
Figure 29: An error example of Wrong Param Type on Normal Data.
{mdframed}
{ "SignalAccuracyEvaluator":{ "device_id":"GPS-7526“ }}
(model output)
{"SignalAccuracyEvaluator":{ "device_id":"GPS-7524“}}
Figure 30: An error example of Wrong Param Value on Normal Data.
{mdframed}
{"EcoDiversityMonitor_assessVegetation": {
    "area_identifier": "Amazon rainforest",
    "detailed_report": true}}
(model output)
Figure 31: An error example of Wrong Output Format on Normal Data.
{mdframed}
(model output)
  "Get_Middle_East_Cities": {
    "sort": "desc",
    "sort_by": "internet_speed",
    "size": 10,
    "page": 1
(The question did not mention the sort method)
Figure 32: An error example of Error Detection on Special Data.
{mdframed}
(model output)
(Find missing necessary parameters, but not correctly identified)
Figure 33: An error example of Error Correction on Special Data.